OpenStudy (anonymous):

Simplify: square root -8 times square root -2

OpenStudy (anonymous):

-4.

OpenStudy (saifoo.khan):

\[\sqrt{-8{\sqrt{-2}}}\] this?

OpenStudy (anonymous):

but i thought you couldn't square root a negative number? The second one looks better

OpenStudy (anonymous):

I got -4. It is i square root 8 minus i square root 2. Which gives you i^2 square root 16. Which is (-1)(4). And gives you the answer -4.

OpenStudy (anonymous):

Shouldn't it be \[\sqrt{-8}\sqrt{-2} = \sqrt{16}?\]

OpenStudy (anonymous):

thats what i was thinking at first... i wasnt sure

OpenStudy (anonymous):

By the way, you will learn later on that you can take the square of a negative number.

OpenStudy (anonymous):

ans \[2^{7/4}\times i ^{3/2}\]

OpenStudy (anonymous):

i is imaginari number its value(\[\sqrt{-1}\])